OpenStudy (anonymous):

solve y=-2x^2 for y=mx+b

OpenStudy (anonymous):

that actually can't be done. y=mx+b is a basic equation for a straight line, but y=-2x^2 is a parabola (more specifically, it opens downward and is skinnier than the basic y=x^2 function)

OpenStudy (anonymous):

so as this bing a problem in a text book, the answer would be no solution?

OpenStudy (anonymous):

correct. y=mx+b is an equation of a straight line, and there is no way to model a parabola as a straight line. if you go to an online graphing calculator, you can type in the y=-2x^2 function and see that it isn't a straight line. =]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem =]